what is p(x<84), when mean is 100 and standard deviation is 8

Answer by stanbon(75887)

what is p(x<84), when mean is 100 and standard deviation is 8
---
z(84) = (84-100)/8 = -16/8 = -2
---
P(x < 84) = P(z < -2) = normalcdf(-100,-2) = 0.0228
=====================
Cheers,
Stan H.